Use the following to answer questions 31-33. . Over the past six months, Seven Flags conducted a marketing study on how to park experience. The study cost $3 million and the results suggested that kids only roller coaster The equipment for the new roller coaster will cost $30 million install improve their ags add a · S5 million to he equipment will be depreciated st and sold for that amount in 20 years. The marketing team at Seven Flags expects attendance at the park to increase by isiOS per year. The average visitor pays $40 per ticket. The operating cost Wi e will go up by $600,000 per year for the increased visitors. The company will increase net working cap This working capital will be liquidated at the end of tl The marginal tax rate is 40%. The cost of capital is 12.00%. ital by $200.000 at the beginning of the project. e project 31. Based on this information, the initial net cash flow of the project (i.e. CHO) is S a. -34,400,000 b. 34,800,000 c -35,200,000 d. -35,600,000 e. -37,000,000 32. Based on this information, the projects total cash flow in year 20 is S a. 11,080,000 b. 11,150,000 c. 11,220,000 d. 11,290,000 11,360,000 53 Bascdon otis odotion the prjetes NPV is S b. 2,939,340.94 4,201,689.48 d. 6,547,312.72 e. 8.586,240.32

Show your work
Answer #1

Question 31) Answer (e)

The initial net cash flow at time=0 (i.e. start of first year) = -(Initial cost + Increase in working capital)

= -(35,000,000 + 200,000) = -37,000,000

Question 32) Answer (c) $11,220,000

Total cash flows for year 20 = Incremental Profit After Tax + Depreciation + Cash flow from Salvage Value + Cash flow from liquidation of working capital

Incremental revenue during year 20 = Incremental revenue every year

Incremental revenue every year = Incremental visitors * price per ticket = 200,000 * 40 = 8,000,000

Incremental costs every year = 600,000

Incremental Earnings Before Depreciation and Tax = 8,000,000 - 600,000 = 7,400,000

Depreciation per year = (Total cost capitalized - Salvage value ) / number of life years

Depreciation per year = (30 + 5 - 6) / 20 = 1.45 million

Incremental profit before tax = EBDT - Dep = 7,400,000 - 1,450,000 = 5,950,000

Tax = 40% * 5,950,000 = 2,380,000

Incremental profit after tax = 5,950,000 - 2,380,000 = 3,570,000

Incremental cash flow per year = PAT + Non-cash expenses (Depreciation) = 3,570,000 + 1,450,000 = 5,020,000

During year 20, there are additional cash flows of (1) cash inflow due to sale of asset equal to salvage value and (2) cash inflow due to liquidation of working capital

(Note: Since asset is sold at book value at year 20, there will be no profit or loss recognized in P&L account. Similarly, the proceeds from sale of asset will not impact P&L account as it is capital cash flow. Liquidation of working capital is assumed at the same cost at which invested)

Total Project cash flow at year 20 = 5,020,000 + 6,000,000 + 200,000 = $11,220,000

Question 33: Answer: $1,139,340.94 (none of the choices match with correct answer)

Total Present Value = Incremental CF per year * PVIFA(12%,19) + 11,220,000 / (1+12%)^20

Total PV = 5,020,000 * ((1+12%)^19-1)/(12%*(1+12%)^19) + 11,220,000 / (1+12%)^20

Total PV = 36,976,199.83 + 116,3141.10 = 38,139,340.94

NPV = Total PV - Total Investment at time=0

NPV = 38,139,341 - 37,000,000 = $1,139,340.94

Correct Answer is NPV = $1,139,340.94 (none of the choices match with correct answer)